Please help me

IP A potter's wheel of radius 6.5 cm rotates with a period of 0.70 s .

Part A

What is the linear speed of a small lump of clay on the rim of the wheel?

Part B

What is the centripetal acceleration of a small lump of clay on the rim of the wheel?

Part C

What is the linear speed if the period of rotation is doubled?

Part D

What is the centripetal acceleration if the period of rotation is doubled?
Answer #1

Part A & B

Radius of the wheel = R = 6.5 cm = 0.065 m

Time period of rotation = T = 0.7 sec

Angular speed of the wheel = \omega

\omega = \frac{2\pi }{T}

\omega = \frac{2\pi }{0.7}

\omega = 8.976 rad/s

Linear speed of a lump of clay on the rim of the wheel = V

V = \omega R

V = (8.976)(0.065)

V = 0.583 m/s

Centripetal acceleration of the lump of clay = a

a = \omega ^{2}R

a = (8.976)^{2}(0.065)

a = 5.24 m/s2

Part C & D

Time period of rotation is doubled.

Radius of the wheel = R = 6.5 cm = 0.065 m

Time period of rotation = T = 2 x 0.7 sec = 1.4 sec

Angular speed of the wheel = \omega

\omega = \frac{2\pi }{T}

\omega = \frac{2\pi }{1.4}

\omega = 4.488 rad/s

Linear speed of a lump of clay on the rim of the wheel = V

V = \omega R

V = (4.488)(0.065)

V = 0.292 m/s

Centripetal acceleration of the lump of clay = a

a = \omega ^{2}R

a = (4.488)^{2}(0.065)

a = 1.31 m/s2

A) Linear speed of a lump of clay on the rim of the wheel = 0.583 m/s

B) Centripetal acceleration of the lump of clay = 5.24 m/s2

C) Linear speed if the period of rotation is doubled = 0.292 m/s

D) Centripetal acceleration if the period of rotation is doubled = 1.31 m/s2